OpenStudy (anonymous):

what are all possible solutions for tan theta= 1

OpenStudy (anonymous):

\[\huge \tan \theta=a \implies \theta=\pi+\arctan a\]

OpenStudy (anonymous):

\[\huge \pi+\arctan a+2\pi k,\forall k \in Z\]

OpenStudy (anonymous):

your a=1

OpenStudy (anonymous):

wait, i still dont get it..